OpenStudy (anonymous):

Joan is running in a straight line such that her position varies with time as f(t) = 2t + 5 where t is the elapsed time (in seconds) and f(t) is the distance from the starting point (in meters). What is the inverse function of f(t)?

OpenStudy (calculusfunctions):

To find the inverse of a function i). rewrite y = f(x) ii). interchange x and y iii). solve for y Example: Find the inverse of f(x) = mx +b I). y = mx + b ii). x = my + b iii). x - b = my y = (x - b)/m Therefore\[f ^{-1}(x)=\frac{ x -b }{ m }\]

OpenStudy (calculusfunctions):

Good Luck!! I have to log out now but I hope that helps!

OpenStudy (anonymous):

I need help :(